OpenStudy (biokiiller396):

If I initially have 4.0 L of gas at a pressure of 1.1 atm, what will the volume be if I increase the pressure to 3.4? (Combined Gas Law)

OpenStudy (accessdenied):

Combined gas law: \( \dfrac{P_1 \times V_1}{T_1} = \dfrac{P_2 \times V_2}{T_2} \) We are not given a temperature, so we should assume it remains constant, and so it can be omitted from the equation (multiply both sides by the same T) and we get Boyle's Law statement: \( P_1 V_1 = P_2 V_2 \) You have an initial volume and initial pressure. Those can be placed in for V_1 and P_1, respectively. You also have the final pressure, so you can plug that in for P_2. That just leaves V_2, right? \( P_1 V_1 = P_2 \color{blue}{V_2} \) <-- with everything else in place, we just solve for V_2

OpenStudy (biokiiller396):

Thank you so much :)

OpenStudy (accessdenied):

Glad to help!